# Flexible MIMO for Future Wireless Communications: Which Flexibilities are Possible?

### v1

@@ Abstract
To enable next-generation wireless communication networks with modest spectrum availability, multiple-input multiple-output (MIMO) technology needs to undergo further evolution. In this paper, we introduce a promising next-generation wireless communication concept: flexible MIMO technology. This technology represents a MIMO technology with flexible physical configurations and integrated applications. We categorize twelve representative flexible MIMO technologies into three major classifications: flexible deployment characteristics-based, flexible geometry characteristics-based, and flexible real-time modifications-based. Then, we provide a comprehensive overview of their fundamental characteristics, potential, and challenges. Furthermore, we demonstrate three vital enablers for the flexible MIMO technology, including efficient channel state information (CSI) acquisition schemes, low-complexity beamforming design, and explainable artificial intelligence (AI)-enabled optimization. Within these areas, eight critical sub-enabling technologies are discussed in detail. Finally, we present two case studies-pre-optimized irregular arrays and cell-free movable antennas-where significant potential for flexible MIMO technologies to enhance the system capacity is showcased.
@@ I Introduction
Multiple-input multiple-output (MIMO) technology has been instrumental in the evolution of wireless networks. It has achieved remarkable success in the fifth-generation (5G) wireless communication networks, relying on its ability to boost spectral efficiency (SE) through spatial multiplexing and beamforming. Looking ahead to future generations, wireless communication networks will face even more stringent requirements across an ever-wider area of application scenarios. To facilitate these requirements and application scenarios, MIMO technology needs to evolve further. In traditional MIMO technology, the base station (BS) is equipped with a uniform planar array (UPA) or a uniform linear array (ULA), which remains static after its deployment. However, this fixed array configuration lacks flexibility and adaptability to face dynamically varying wireless channel environments and enable future application scenarios, such as immersive communication. Thus, it is essential to explore and develop next?generation MIMO technologies that offer geometric and topological reconfiguration, empowering future wireless systems with the performance and versatility they require [1, 2].
Motivated by the above observations, in this paper, we propose a new overarching concept: flexible MIMO technology. Flexible MIMO denotes a MIMO technology with flexible physical configurations and integrated applications, including flexible deployment characteristics, geometry characteristics, and real-time modified configurations. In flexible MIMO technology, the network topology, integrated application, matrix geometry characteristic, and real-time antenna movement can be configured flexibly and dynamically. More specifically, the network topology denotes the network device and the network deployment strategy. The integrated application represents the services enabled by the MIMO technology. The characteristic of the array geometry includes the shape of the array, the orientation/brightness of the array, and the antenna deployment strategy. Real-time antenna movement is the further modification strategy beyond the above technologies, empowered by instantaneous channel state information (CSI). Indeed, by flexibly reconfiguring multi-antenna arrays in the above four major aspects, the flexible MIMO technology can be effectively utilized to empower future wireless communication networks.
@@
In this paper, we investigate a novel concept: flexible MIMO technology to empower future wireless communication networks. The major contributions are summarized as follows.
  • We classify and introduce twelve different flexible MIMO technologies, organized into three categories: flexible deployment characteristics based on, flexible geometry characteristics based on, and flexible real-time modifications based on. Then, we introduce the fundamentals of these technologies and summarize promising research clues.
@@
  • In case studies, we study two promising flexible MIMO technologies, pre-optimized irregular arrays, and cell-free movable antennas. Then, we demonstrate the promising potential for flexible MIMO technologies to enhance the network capacity.
@@ II-A Flexible Deployment Characteristics-Based Technology
We will now introduce promising flexible MIMO technologies based on the above flexibility sources. The first classification is the flexible deployment characteristics-based technology, which can be further classified from the network topology and integrated application-tailored one.
@@ II-A1 Flexible Network Topology-Based
Recently, a network topology called the PASS is emerging [4]. In PASS, a dielectric waveguide with low propagation loss is applied, where small separated dielectrics, typically attached to the tips of plastic clips or clothespins, are pinched in the waveguide. Compared with leaky feeder cable, typically applied in tunnels, the number and positions of pinched dielectrics can be flexibly adjusted by pinching them in or detaching them from the waveguide. By incorporating several waveguides in the BS, PASS can be flexibly utilized, where the near-wired transmission can be achieved and the stable line-of-sight (LoS) propagation can be established between dielectrics and UEs.
@@ II-A2 Flexible Application-Tailored
TABLE I: Comparison of all studied flexible MIMO technologies.
@@ II-B Flexible Geometry Characteristics-Based Technology
The conventional uniform array may not always be the best topology choice. Instead, designing non-uniform or irregular arrays, exploiting the flexibility of antenna arrangement, can enhance the system performance compared to the uniform array. The array can achieve the desired spatial resolution in the directions that matter, and sacrifice other directions. Indeed, a pre-optimized irregular array can be derived, where antenna elements can be arbitrarily deployed within predefined non-overlapping rectangular areas [9]. The antenna arrangement strategy can be “pre-optimized” before the deployment of the array on-site based on the statistical information with respect to numerous possible user distributions in a particular coverage area. Different from the instantaneous information-based MA technology [5] introduced in the next part, this pre-optimized irregular array can be designed based on statistical information with low complexity but can sometimes achieve about $90\%$ performance for the MA technology. Notably, irregular arrays are particularly useful when the arrays are sparse, where there is room for irregular arrangement. Compared with uniform sparse arrays, there are no grating lobes in irregular arrays, even if there will be large side-lobes.
@@ II-C Flexible Real-Time Modifications-Based Technology
The previously discussed MIMO technologies feature flexible deployment and geometry, but are either optimized once and for all, or modified slowly compared to the channel coherence time. In this part, we discuss flexible MIMO technologies leveraging real-time modifications, driven by instantaneous small-scale channel fading properties at a pace faster than the coherence time (i.e., at the millisecond level). Note that some of the previously introduced technologies, such as rotatable array technology, can also be adapted based on instantaneous channel fading characteristics. However, for clarity in differentiating various flexibility features, we classify these flexible MIMO technologies as shown in Fig. 1. Specifically, these real-time modification¨Cbased technologies rely on antenna movements driven by instantaneous channel fading properties, which can be further categorized into movements along the array and movements perpendicular to the array.
@@
In the previous technology, antenna movements are along the array. However, with the aid of micro- and nano-fabrication and flexible metamaterials, the antenna elements can move perpendicular to the array to showcase soft characteristics in the perspective of the whole array. This technology was first studied in the field of EM and the authors in [13] applied this technology to wireless communication networks and called it as flexible intelligent metasurfaces (FIMs). Specifically, each antenna element can flexibly move perpendicular to the array within a confined region with the aid of a controller. By doing this, the array can morph its 3D shape in a soft feature. The angles of arrival (AoA) and angles of departure (AoD) can be dynamically adjusted and thus the channel response also varies as the the antenna movement. Indeed, the antenna movement strategy perpendicular to the array can be optimized dynamically based on the instantaneous channel information. In summary, we compare all studied technologies in Table I. Meanwhile, we motivate the design ideas and summarize some insightful research clues as in Fig. 2.
@@ III-A3 Channel Map
The channel map or channel charting is an environment-aware framework, tagged with the locations of transceivers, which is a novel concept in digital twins and can provide location-specific CSI based on inputted geographical locations. The channel map works as a site-specific dataset equipped at stations, which is expected to generate the desired CSI, such as the channel gain, path loss, and even the real-time small-scale fading channel coefficients. Note that the channel map is environment-empowered, which can be efficiently targeted at the specific area. This novel methodology can be efficiently incorporated into flexible MIMO technologies. For instance, relying on the abundant historical channel measurements and accurate radio propagation depiction? of the channel map, we can access the potential user distributions and propagation environment features to efficiently pre-design the flexible multi-antenna arrays, such as T9 and T11. More importantly, the channel map methodology can also empower the above three CSI acquisition methodologies with the aid of its prior environment-enabled framework.
@@ III-B1 Statistical Information-Based Design
Compared with rapidly changing small-scale instantaneous information, statistical information remains constant for a long period of time. Thus, it would be promising to implement the statistical information-based design to lower the computational complexity compared to the conventional instantaneous information-based one. [14] proposes statistical information-cored receive combining schemes, where statistics-based matrices replace the instantaneous information-based matrix inversion part in MMSE combining, which significantly reduces the computational complexity. It can be observed that approaching SE performance to that of optimal MMSE combining can be achieved by this beamforming scheme for scenarios with low mobility users and severe pilot contamination. Focusing on future flexible MIMO technologies, it is also vital to implement CSI-based MMSE-like combining/precoding in most practical scenarios.
@@ III-C1 Array Configuration Optimization Design
As discussed above, the geometry characteristics and antenna movement for the array can be flexibly adjusted to provide enhanced performance. The adjustment strategy design can be viewed as the optimized problem. In practice, however, directly solving such optimization problems can be computationally demanding. Therefore, the introduction of XAI-based optimization methods into array configuration optimization emerges as a promising solution. Specifically, by exploiting XAI¡¯s transparent optimization pathways, the inherent complexity of determining optimal array geometry characteristics and antenna movement strategies can be effectively managed. For instance, an XAI model utilizing Bayesian inference can quantify uncertainties associated with antenna placement decisions, thereby guiding robust and interpretable optimization strategies.
@@ III-C2 Beamforming Optimization Design
For the beamforming optimization design, the XAI model, leveraging deep unfolding techniques, can enhance the adaptability and interpretability of beamforming solutions. Deep unfolding can explicitly construct the optimization process as iterative and interpretable steps, where the gradient descent method is employed to iteratively refine the beamforming scheme based on local channel feedback. Through this structured and transparent optimization pathway, the XAI model effectively balances the desired signals and interference, enabling precise adjustment of the beamforming scheme. Consequently, the resulting beamforming scheme exhibits improved transmission efficiency and robustness, particularly under stringent resource constraints and dynamic channel environments.
@@ IV Case Studies
IV-A Pre-Optimized Irregular Array
In this part, we bring a numerical case study for the pre-optimized irregular array (PIA) technology introduced in T9. As introduced in T9, the PIA can be designed based on numerous possible user distributions within a specific area. We also include the following comparison benchmarks: MA-enabled array, where the array configurations are optimized based on each realization of the user distributions, uniform sparse planar array (USPA) with the BS equipped with a UPA with antenna spacing being $5\lambda$, and half-wavelength planar array (HwPA) with the BS equipped with a UPA with the half-wavelength antenna spacing. Figure 4 showcases the cumulative distribution function (CDF) of the sum rate for the PIA scheme and the other benchmark schemes. As observed, the PIA can significantly outperform other fixed array configurations, and there is only a $6\%$ gap for the average sum rate between the PIA and MA-enabled array. This observation provides useful insights for the future flexible MIMO technology.
Figure 4: The CDF of the sum rate for different array schemes with one BS equipped with $6\times 6$ antennas and $8$ users with $2$ antennas each. The location of each BS antenna can be adjusted within a local two-dimensional (2D) region of size $5\lambda\times 5\lambda$ for both the PIA and MA technologies. The optimization of antenna locations in PIA and MA technologies can be implemented based on the particle swarm optimization (PSO) algorithm. The number of realizations for user distributions is $1000$. The other simulation parameters follow from those of [9] and more technical results can be observed in [9].
Figure 5: Sum SE for different architectures versus the signal-to-noise ratio (SNR) with four APs and $4$ users. All APs are equipped with UPAs with $2\times 2$ antennas each, and all users are equipped with a single antenna. Antennas at each AP can flexibly move along the array. The multi-path array response-based channel model is utilized and the carrier frequency is $3.1\,\mathrm{GHz}$. The terminologies “Centralized” and “Distributed” denote the centralized and distributed optimization frameworks, respectively. The local minimum mean-square error (LMMSE) precoding denotes the distributed precoding scheme at each AP [14].

### v2

@@ Abstract
In conventional multiple-input multiple-output (MIMO), static array configurations struggle in dynamic environments, and further antenna scaling is bounded by cost, energy, and footprint. Emerging approaches, which can enable next-generation wireless communication networks with modest spectrum availability by leveraging flexibility and adaptability rather than sheer array growth, are therefore needed. In this paper, we present a taxonomy framework, referred to as flexible MIMO technology, that systematically categorizes a wide range of evolving MIMO technologies. The focus is on MIMO technologies with flexible physical configurations and integrated applications. We categorize twelve representative flexible MIMO technologies into three major classifications: flexible deployment characteristics-based, flexible geometry characteristics-based, and flexible real-time modifications-based. We then comprehensively overview their fundamental characteristics, potential, and challenges. In addition, we highlight three vital enablers for flexible MIMO technology, including efficient channel state information acquisition schemes, low-complexity beamforming design, and explainable artificial intelligence (AI)-enabled optimization, and discuss eight representative sub-techniques. Finally, two brief case studies—pre-optimized irregular array for high-speed railway network and cell-free movable antenna—are presented, showing how flexible MIMO can open new design possibilities and inspire future research directions for next-generation wireless networks.
@@ I Introduction
[…]
Motivated by the above observations, in this paper, we propose a new overarching taxonomy framework: flexible MIMO technology. Flexible MIMO denotes a MIMO technology with flexible physical configurations and integrated applications, including flexible deployment characteristics, geometry characteristics, and real-time modified configurations. In flexible MIMO technology, the network topology, integrated application, matrix geometry characteristic, and real-time antenna movement can be configured flexibly and dynamically. More specifically, the network topology denotes the network device and the network deployment strategy. The integrated application represents the services enabled by the MIMO technology. The characteristic of the array geometry includes the shape of the array, the orientation/brightness of the array, and the antenna deployment strategy. Real-time antenna movement is the further modification strategy beyond the above technologies, empowered by instantaneous channel state information (CSI). Indeed, by flexibly reconfiguring multi-antenna arrays in the above four major aspects, the flexible MIMO technology can be effectively utilized to empower future wireless communication networks. The purpose of introducing “flexible MIMO” is not to coin another isolated label, but to establish a taxonomy framework for many emerging MIMO technologies. While numerous concepts have appeared, they are often presented separately and lack a holistic vision. This framework, motivated by the new MIMO flexibilities, enables wireless practitioners to adopt a generalized perspective and derive clear, systematic research clues for the future evolution of MIMO.
@@
In this paper, we investigate a novel taxonomy framework: flexible MIMO technology to empower future wireless communication networks. The major contributions are summarized as follows.
  • We classify and introduce twelve flexible MIMO technologies, organized into three categories: flexible deployment characteristics based on, flexible geometry characteristics based on, and flexible real-time modifications based on. Then, we introduce the fundamentals of these technologies and bring promising design motivations.
@@
  • In case studies, we study two promising flexible MIMO technologies, pre-optimized irregular arrays for high-speed railway networks, and cell-free movable antennas. Then, we demonstrate the promising potential for flexible MIMO technologies to enhance the network capacity.
@@ II-A Flexible Deployment Characteristics-Based Technology
We will now introduce promising flexible MIMO technologies based on the above flexibility sources as shown in Fig. 1. The first classification is the flexible deployment characteristics-based technology, which can be further classified from the network topology and integrated application-tailored one.
@@ II-A1 Flexible Network Topology-Based
A recent widely studied network topology, the PASS [4], employs a low-loss dielectric waveguide with clip-on separated dielectrics (“pinching antennas”). Unlike leaky-feeder cables used in tunnels, PASS allows the number and positions of pinched dielectrics to be adjusted by simple attach/detach operations across the waveguide. By incorporating multiple waveguides at the BS, PASS enables near-wired transmission and stable line-of-sight (LoS) links between the dielectrics and UEs. The insertion loss along the waveguide can be described using models such as [4, Eq. (2)]. Considering the linear feature of the waveguide, the practical outdoor deployment of PASS is one important open challenge.
@@ II-A2 Flexible Application-Tailored
TABLE I: Characteristics, potentials, and key challenges of all studied flexible MIMO technologies.
@@ II-B Flexible Geometry Characteristics-Based Technology
The conventional uniform array may not always be the best topology choice. Instead, designing non-uniform or irregular arrays, exploiting the flexibility of antenna arrangement, can enhance the system performance compared to the uniform array. The array can achieve the desired spatial resolution in the directions that matter, and sacrifice other directions. Indeed, a pre-optimized irregular array can be derived, where antenna elements can be arbitrarily deployed within predefined non-overlapping rectangular areas [9]. The antenna arrangement strategy can be “pre-optimized” before the deployment of the array on-site based on the statistical information with respect to numerous possible user distributions in a particular coverage area. Different from the instantaneous information-based MA technology [5] introduced in the next part, this pre-optimized irregular array can be designed based on statistical information with low-complexity but can sometimes achieve about $90\%$ performance for the MA technology. Notably, irregular arrays are particularly useful when the arrays are sparse, where there is room for irregular arrangement. Compared with uniform sparse arrays, there are no grating lobes in irregular arrays, even if there will be large side-lobes.
@@ II-C Flexible Real-Time Modifications-Based Technology
The previously discussed MIMO technologies feature flexible deployment and geometry, but are either optimized once and for all, or modified slowly compared to the channel coherence time. In this part, we discuss flexible MIMO technologies leveraging real-time modifications, driven by instantaneous small-scale channel fading properties at a pace faster than the coherence time (i.e., at the millisecond level). Specifically, these real-time modification–based technologies rely on antenna movements driven by instantaneous channel fading properties, which can be further categorized into movements along the array and movements perpendicular to the array. Note that some of the previously introduced technologies, such as rotatable array technology, can also be adapted based on instantaneous channel fading characteristics. The specific implementation rule can be flexibly chosen based on practical requirements.
@@
In the previous technology, antenna movements are along the array. However, with the aid of micro- and nano-fabrication and flexible metamaterials, the antenna elements can move perpendicular to the array to showcase soft characteristics in the perspective of the whole array. This technology was first studied in the field of EM and the authors in [13] applied this technology to wireless communication networks and called it as flexible intelligent metasurfaces (FIMs). Specifically, each antenna element can flexibly move perpendicular to the array within a confined region with the aid of a controller. By doing this, the array can morph its 3D shape in a soft feature. The angles of arrival (AoA) and angles of departure (AoD) can be dynamically adjusted and thus the channel response also varies as the the antenna movement. Indeed, the antenna movement strategy perpendicular to the array can be optimized dynamically based on the instantaneous channel information.
II-D Lessons Learned
All these twelve technologies have the potential to be deployed in future wireless networks, which have respective prototypes in different developing stages. Some technologies, such as T1 and T2, are already mature enough to be implemented in practical communication scenarios. Some other technologies, such as T4, T11, and T12, are still in the early stages with some initial prototypes. Beyond these twelve technologies, electronically reconfigurable architectures, such as electronically steerable parasitic array radiators (ESPARs), reconfigurable-pixel antennas, and dynamic metasurface antennas (DMAs), have also been extensively studied. Compared with mechanically reconfigurable systems (e.g., movable/fluid antennas), these electronically reconfigurable architectures achieve compactness, low latency, and high integration capabilities, making them mature and well-suited for near-term deployment. Mechanical approaches provide greater adaptability but depend on motors or actuators, likely resulting in higher cost, power consumption, and slower reconfiguration. Viewed together, electronic and mechanical reconfigurability are complementary: electronic solutions offer practical deployment today, while mechanical solutions explore the ultimate limits of flexibility and inspire future designs, where they define a broader and more balanced landscape for the evolution of flexible MIMO technologies. While mechanical reconfigurability provides valuable insights into flexible MIMO, its scalability and practicality remain an open challenge.
Note that the aim of introducing the terminology “flexible MIMO” is not to coin another isolated label or rebrand the existing MIMO technologies, but to formulate a useful taxonomy framework for many promising MIMO technologies. “Flexibility” is an essential property that connects many emerging MIMO technologies, where different flexibilities for MIMO can be explored to formulate an integrated MIMO design framework. For potential readers and practitioners, we want to share the following motivations.
First, we motivate the design ideas and provide some insightful research clues as in Fig. 2. One intuitive implementation of future flexible MIMO technology can be motivated based on different requirements. One needs to determine “What network topology should be used?” based on the fundamentals in T1-T4. Then, to explore “What integrated application is tailored for?”, one can capture the useful insights from T5-T7. Moreover, relying on the focused network topology and integrated applications, one can further consider “What array geometry feature should be selected?” from the array level based on T8-T10. Finally, one can decide “Should further real-time modification be applied?” to implement the instantaneous information-based modification as in T11-T12.
Second, since twelve emerging MIMO technologies that provide different kinds of flexibility are studied in this paper, to also facilitate the actionable depth of these technologies, we summarize the major aspects of all of them in Table I. More specifically, we describe the main technical insights for all flexible MIMO technologies, such as the major ideas, representative characteristics, featured capabilities, intractable technical problems, and practical implementation challenges etc., through Table I.
Third, we stress that providing a unified quantitative comparison across all twelve flexible MIMO technologies is not feasible at this stage, as they aim to address different challenges, and it is also not our intention. Many technologies are still in early exploration, with heterogeneous hardware assumptions and distinct operating scenarios, which makes a fair and comprehensive evaluation impossible at this stage. Instead, we aim to provide a taxonomy for these technologies with the clue of “flexibility” and highlight key characteristics in Table I, design guidelines in Fig. 2. We hope MIMO technology can be significantly evolved in all these potential directions.
Fourth, the taxonomy framework, relying on three major classifications, does not mean that each MIMO technology should be totally restricted to one particular classification, where a practical system can make use of multiple of these classes. For instance, rotatable array technology can also be implemented based on the real-time instantaneous channel. However, these observations are not a drawback of our taxonomy framework, but showcase that these technologies can be flexibly utilized based on different practical requirements. We also advocate for utilizing the studied MIMO technologies in this paper with a synergistic and flexible vision, instead of limiting one particular technology to one particular classification. Specifically, we provide two case studies that synergistically combine some of the listed technologies promisingly.
@@ III-A3 Channel Map
The channel map or channel charting is an environment-aware framework, tagged with the locations of transceivers, which is a novel concept in digital twins and can provide location-specific CSI based on inputted geographical locations. The channel map works as a site-specific dataset equipped at stations, which is expected to generate the desired CSI, such as the channel gain, path loss, and even the real-time small-scale fading channel coefficients. Note that the channel map is environment-empowered, which can be efficiently targeted at the specific area. This novel methodology can be efficiently incorporated into flexible MIMO technologies. For instance, relying on the abundant historical channel measurements and accurate radio propagation depiction of the channel map, we can access the potential user distributions and propagation environment features to efficiently pre-design the flexible multi-antenna arrays, such as T9 and T11. More importantly, the channel map methodology can also empower the above three CSI acquisition methodologies with the aid of its prior environment-enabled framework.
One important electromagnetic aspect is mutual coupling, which becomes particularly critical in mechanically reconfigurable architectures. Notably, one main reason to consider sparse arrays (fixed or with movable antennas) is that the mutual coupling becomes small. As antenna positions or orientations change, the mutual coupling among elements varies, altering the effective channels. To efficiently acquire CSI, the joint estimation scheme that can dynamically disentangle the propagation channel response from the coupling effects can be viewed as an important method. This remains a key open research problem at the intersection of electromagnetic theory and channel estimation.
@@ III-B1 Statistical Information-Based Design
Compared with rapidly changing small-scale instantaneous information, statistical information remains constant for a long period of time. Thus, it would be promising to implement the statistical information-based design to lower the computational complexity compared to the conventional instantaneous information-based one. [14] proposes statistical information-cored receive combining schemes, where statistics-based matrices replace the instantaneous information-based matrix inversion part in MMSE combining. It can be observed that approaching SE performance to that of optimal MMSE combining can be achieved by this beamforming scheme for scenarios with low mobility users and severe pilot contamination. This methodology involves a higher offline precomputation load than classical MMSE combining. However, because it relies on statistical information, the overhead is infrequent and acceptable.
@@ III-C1 Array Configuration Optimization Design
As discussed above, the geometry characteristics and antenna movement for the array can be flexibly adjusted to provide enhanced performance. The adjustment strategy design can be viewed as the optimized problem. In practice, however, directly solving such optimization problems can be computationally demanding. Therefore, the introduction of XAI-based optimization methods into array configuration optimization emerges as a promising solution. Specifically, by exploiting XAI’s transparent optimization pathways, the inherent complexity of determining optimal array geometry characteristics and antenna movement strategies can be effectively managed. For instance, an XAI model utilizing Bayesian inference can quantify uncertainties associated with antenna placement decisions, thereby guiding robust and interpretable optimization strategies.
@@ III-C2 Beamforming Optimization Design
For the beamforming optimization design, the XAI model, leveraging deep unfolding techniques, can enhance the adaptability and interpretability of beamforming solutions. Deep unfolding can explicitly construct the optimization process as iterative and interpretable steps, where the gradient descent method is employed to iteratively refine the beamforming scheme based on local channel feedback. Through this structured and transparent optimization pathway, the XAI model effectively balances the desired signals and interference, enabling precise adjustment of the beamforming scheme. Consequently, the resulting beamforming scheme exhibits improved transmission efficiency and robustness, particularly under stringent resource constraints and dynamic channel environments. However, XAI entails nontrivial development and deployment costs. Training is often compute- and data-intensive, requiring large, high-quality datasets and incurring substantial training overhead in time and energy. Moreover, the online complexity of some deep-unfolded or white-box models can exceed that of conventional iterative algorithms, underscoring a trade-off among explainability, performance, and real-time feasibility that must be assessed for the target application.
@@ IV Case Studies
Figure 4: Average SE against the BS coverage area radius for the HSR scenario with $350\,\mathrm{km/h}$ speed. The coverage area radius denotes the distance between the maximum coverage distance and the center of the BS. The BS is equipped with $6\times 6$ antennas, serving 88 single-antenna UEs (carriages) with $28\,\mathrm{m}$ relative height. The location of each BS antenna can be adjusted within a local two-dimensional (2D) region of size $3\lambda\times 3\lambda$ for both the PIA and MA technologies. We sample 8080 realizations of potential positions of carriages along the practical measured railway operation trajectory to optimize the PIA.
IV-A Pre-Optimized Irregular Array-Enabled HSR
In this part, we utilize the pre-optimized irregular array (PIA) technology introduced in T9 in the HSR communication introduced in T7. As introduced in T9, the PIA can be designed based on statistical knowledge of the user distributions within a specific area, which suits the HSR communication scenario with its specific operating area and predictive movement trajectory. We also include the following comparison benchmarks: MA-enabled array, where the array configurations are real-time optimized based on each realization of the user distributions, uniform sparse planar array (USPA) with the BS equipped with a UPA with antenna spacing being $3\lambda$, and half-wavelength planar array (HwPA) with the BS equipped with a UPA with the half-wavelength antenna spacing. Figure 4 showcases the average SE against the BS coverage area radius for the PIA and the other benchmark schemes. We observe that the PIA outperforms other fixed array configurations, and there is only about $2\%$ gap in average SE between the PIA and the MA-enabled array in the considered setup. This observation demonstrates the potential of synergistically integrating different flexible MIMO technologies.
Figure 5: Sum SE for different architectures versus the signal-to-noise ratio (SNR) with 44 APs and 44 users. All APs are equipped with UPAs with $2\times 2$ isotropic antennas each, and all users are equipped with a single isotropic antenna with $25\,\mathrm{m}$ spacing. The multi-path array response-based channel model without mutual coupling is utilized, and the carrier frequency is $3.1\,\mathrm{GHz}$. The terminologies “Centralized” and “Distributed” denote the centralized and distributed optimization frameworks, respectively.